A 16 -inch laptop has a screen that is 7 inches tall. How wide is the screen?

Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)


A 16 -inch laptop has a screen that is 7 inches tall
diagonal of the rectangular TV =16 "
Height = 7 "
width (w) = ?
We apply Pythagoras theorem
w^2+h^2= 16^2
w^2 = 16^2-(7^2)
w^2= 207
w = sqrt(207)
Simplify

Answer by Alan3354(69443)

A 16 -inch laptop has a screen that is 7 inches tall. How wide is the screen?
----------
If the diagonal is 16 inches:
w^2 + 7^2 = 16^2